OpenStudy (anonymous):

Can someone explain how to get the answer to this problem? Preferably step by step please! If 5^{-x+3}=2, then 5^{x-2}= ? Thank you!! :)

OpenStudy (amistre64):

you want to find a value of x that makes it seem plausible but only when you take a quick glance :)

OpenStudy (amistre64):

(-x+3) ln(5) = ln(2) -x+3 = ln(2)/ln(5) -x = (ln(2)/ln(5)) - 3 x = -(ln(2)/ln(5)) + 3 insert that value into the other side to solve

OpenStudy (anonymous):

I used log instead of ln, it's the same thing right?

OpenStudy (amistre64):

yeah, i used ln casue its quicker to type than log and not as much room for error ;)

OpenStudy (amistre64):

you coulda used log5 as well, but then log5(2) seems superfluous

OpenStudy (anonymous):

For some reason, I'm getting weird answers instead of the choices given. The choices are 10, 25, 5/2 and 2/5.

OpenStudy (amistre64):

for x or for the new value?

OpenStudy (anonymous):

I typed the question exactly the way it's stated. I'm guessing it's after the value?

OpenStudy (amistre64):

5^(-(ln(2)/ln(5)) + 3 - 2) 5^(-(ln(2)/ln(5)) + 1) = 2.5 http://www.wolframalpha.com/input/?i=5^%28-%28ln%282%29%2Fln%285%29%29+%2B+1%29

OpenStudy (amistre64):

2.5 = 2' 1/2 = 5/2

OpenStudy (anonymous):

Wow! Thank you!

OpenStudy (amistre64):

yw :)